# TRIM: Token-wise Attention-Derived Saliency for Data-Efficient Instruction Tuning

### v1

@@ Appendix D Training Datasets Overview
Our selection pool $S$ is the union of four public instruction-tuning sources, Dolly, CoT, Oasst1, and Flan_v2, totaling 270,679 examples after filtering (Table 5). We obtain the processed datasets from LESS repository (Xia et al., 2024). Unless otherwise stated, we do not rebalance per-source proportions prior to selection; methods operate over the pooled corpus $S$ with fixed preprocessing and tokenization settings across all experiments.
@@
Instruction following;
everyday tasks, Q&A,
and summaries.
@@
Chain-of-thought style
prompts and responses.
@@
Multi-turn chat-style
conversations.
@@
Mixture of diverse tasks
for instruction finetuning.
@@
Appendix E Target Datasets & Evaluation Setup
@@
We evaluate on four standard targets, CommonsenseQA, SocialIQA, HellaSwag, and Gsm8K, following the OpenAI-style instruction-tuning evaluation protocol (Wang et al., 2023). For selection, we use a small few-shot validation set $T_{\text{val}}$ of five labeled examples per task (ten for Gsm8K), with no access to test labels. At test time, we use fixed in-context exemplars with the shot counts in Table˜6 (5/0/5/8 for CommonsenseQA/SocialIQA/HellaSwag/Gsm8K), sampled from development pools disjoint from $T_{\text{val}}$. We report accuracy for the multiple-choice tasks and exact match on the final numeric answer for Gsm8K. All prompts and decoding settings are held fixed across methods; no task-specific hyperparameters are tuned on test sets. Datasets are obtained via HuggingFace datasets (Wolf et al., 2020), and evaluation is performed with the EleutherAI lm-evaluation-harness (Sutawika et al., 2023).
Task #Val #Test Eval Shots
CommonsenseQA 5 1,221 5
SocialIQA 5 1,954 0
HellaSwag 5 10,042 5
Gsm8K 10 1,319 8
Table 6: Target datasets and evaluation protocol. “#Val” is the few-shot target validation set $T_{\text{val}}$ used for selection; “Eval Shots” is the number of in-context examples at test time.
Appendix F Training Details and Hyperparameters
F.1 Artifact Licenses
According to their license, all the LLMs used in this paper fall under acceptable use cases. The licenses for the models are linked for perusal: Llama-2-7b, Llama-3.2-1B, Llama-3.1-8b and, Mistral 7B (v0.3).
All software dependencies, including PyTorch and torchvision, are open-source and distributed under MIT or BSD-compatible licenses.
F.2 Implementation Details and Shared Hyperparameters
All models are trained with the same setup, closely following LESS (Xia et al., 2024), using an identical training recipe for data loading, preprocessing, optimizer/schedule, precision, and LoRA configuration; only the underlying architecture varies unless otherwise noted.
@@ F.2 Implementation Details and Shared Hyperparameters
F.3 Model Size and Computational Budget
We evaluate models from 1B to 8B parameters using a shared training/adaptation recipe (Appendix F.2). Concretely, our backbones are Llama-3.2-1B ($\approx$ 1B parameters), Llama-2-7B ($\approx$ 7B parameters), Mistral-7B ($\approx$ 7B parameters), and Llama-3.1-8B ($\approx$ 8B parameters). All runs were executed on a single NVIDIA H200 GPU on an internal cluster (CUDA 12.1, PyTorch 2.1, HuggingFace Transformers 4.46). We do not report wall-clock time or GPU-hours, as these depend on dataset composition and cluster load; instead, we provide asymptotic computational cost and scaling discussion in Section˜5.
Appendix G Ablations on CommonsenseQA (Llama-3.2-1B)
All ablations are run on CommonsenseQA with Llama-3.2-1B at a fixed budget $p=5\%$ and identical fine-tuning hyperparameters. Row entries modify only one factor relative to the default (ours) configuration, which is marked with $\dagger$.
Setting CSQA Acc. (%)
@@ F.3 Model Size and Computational Budget
None 36.27 (0.2)
@@
0.5 39.35 (0.84)
@@
Max only 38.54 (0.4)
Mean only 39.02 (0.7)
@@
Pretrained (no warmup) 37.85 (0.5)
Checkpoint #1 (early) 38.78 (0.3)
@@
Last 3 38.35 (0.6)
@@
Last 12 40.68 (0.3)
Table 8: Ablations on the different hyperparameters that can be used for TRIM for choosing a $5\%$ coreset of CommonsenseQA (CSQA) on a Llama-3.2-1B model. Each section tweaks one factor; and all others stay at the default (marked $\dagger$): OOV penalty =1.0=1.0, pooling =0.5=0.5 mean ++ $0.5$ max, coverage $\eta=0.05$, scoring checkpoint == last, layers == last 6. The experimental results reported in the main paper, use these default values.
Appendix H Subset Distribution of Selected Examples
We present the distribution of selected examples across different data selection methods in Figure 5. The three approaches exhibit substantially different preferences for data from each training subset. However, we observe that there is not a monotonic relationship between the proportion of data selected from a subset and overall method performance. This suggests that each subset contains valuable examples for the target tasks, and the primary challenge lies in identifying the most task-relevant instances.
Notably, TAGCOS, being a task-agnostic method, maintains nearly identical selection distributions across all four tasks (approximately 38.5% CoT, 35.0% Flan_v2, 20.7% Oasst1, and 5.8% Dolly). In contrast, both TRIM and LESS demonstrate task-adaptive behavior, substantially varying their selection patterns based on the specific characteristics of each target task.
In our specific observations, we find that TRIM predominantly selects data from the Flan_v2 dataset for both SocialIQA (81.3%) and CommonsenseQA (71.8%). This preference appears intuitive, as both tasks involve question answering with multiple choice or short answer formats, which are abundantly represented in Flan_v2’s diverse instruction-following examples. For GSM8K, TRIM exhibits a strong shift toward the CoT dataset (67.2%), while reducing Flan_v2 to 32.0%. This selection pattern aligns well with GSM8K’s focus on mathematical reasoning, where step-by-step problem solving, a hallmark of the CoT dataset, is essential. Finally, for Hellaswag, TRIM adopts a more balanced approach, selecting comparable amounts from Flan_v2 (56.0%) and CoT (42.7%). This distribution reflects Hellaswag’s nature as a commonsense inference task that benefits from both natural language understanding examples and reasoning-based training data.

### v2

@@ Appendix E Training Datasets Overview
Our selection pool $S$ is the union of four public instruction-tuning sources, Dolly, CoT, Oasst1, and Flan_v2, totaling 270,679 examples after filtering (Table 6). To unify heterogeneous sources, we format all datasets into a chat-style schema with special tokens (<|user|>, <|assistant|>) and an end-of-text marker, following the instruction-tuning pipeline of Wang et al. (2023). We obtain the processed datasets from LESS repository (Xia et al., 2024). Unless otherwise stated, we do not rebalance per-source proportions prior to selection; methods operate over the pooled corpus $S$ with fixed preprocessing and tokenization settings across all experiments.
@@
Instruction following; everyday tasks, Q&A, and summaries.
@@
Chain-of-thought style prompts and responses.
@@
Multi-turn chat-style conversations.
@@
Mixture of diverse tasks for instruction finetuning.
@@
Appendix F Target Datasets & Evaluation Setup
@@
We evaluate our method across a diverse set of reasoning and question-answering benchmarks. The main paper reports results on MMLU, BBH, TyDiQA, and Gsm8K, while Appendix N additionally includes experiments on CommonsenseQA, SocialIQA, and HellaSwag for completeness and comparison with prior work. It is to be noted that $T_{\text{val}}$ is disjoint from the Test set.
Table 7: Target datasets and evaluation protocol. “#Val” denotes the size of the few-shot target validation set $T_{\text{val}}$ used for data selection; “Eval Shots” denotes the number of in-context examples used at test time. Results for MMLU, BBH, TyDiQA, and Gsm8K are reported in the main paper, while results for CommonsenseQA, SocialIQA, and HellaSwag are included in the appendix.
Dataset #Tasks #Val #Test Eval Shots Answer Type / Task
MMLU 57 285 18,721 5 Multiple-choice (letter)
BBH 27 81 920 3 CoT + final answer
TyDiQA 9 9 1,713 1 Extractive span
CommonsenseQA 1 5 1,221 5 Multiple-choice
SocialIQA 1 5 1,954 0 Multiple-choice
HellaSwag 1 5 10,042 5 Multiple-choice (ending)
Gsm8K 1 10 1,319 8 Numeric answer
All experiments follow a unified OpenAI-style instruction-tuning evaluation protocol (Wang et al., 2023). For data selection, we use a small few-shot target validation set $T_{\text{val}}$ for each (sub)task, with the size determined by the few-shot configuration of the benchmark. Specifically, we use five labeled examples per task for most benchmarks and ten for Gsm8K, with no access to test labels during selection.
At evaluation time, we use fixed in-context exemplars with the shot counts shown in Table 7, sampled from development pools disjoint from $T_{\text{val}}$. We report accuracy for multiple-choice benchmarks, exact match for span-extraction tasks (TyDiQA), and exact match on the final numeric answer for Gsm8K. All prompts, decoding parameters, and evaluation settings are held fixed across methods; no task-specific hyperparameters are tuned on test sets.
Datasets are obtained via the HuggingFace datasets library (Wolf et al., 2020), and evaluation is performed using the EleutherAI lm-evaluation-harness (Sutawika et al., 2023).
Appendix G Training Details and Hyperparameters
G.1 Artifact Licenses
According to their license, all the LLMs used in this paper fall under acceptable use cases. The licenses for the models are linked for perusal: Llama-2-7b, Llama-2-13b, Llama-3.2-1B, Llama-3.1-8b, Pythia-70M and, Mistral 7B (v0.3). All software dependencies, including PyTorch and torchvision, are open-source and distributed under MIT or BSD-compatible licenses.
G.2 Implementation Details and Shared Hyperparameters
All models in Sections 4.1, 4.2, 4.3, and Section 4.4 are trained with the same setup, closely following LESS (Xia et al., 2024), using an identical training recipe for data loading, preprocessing, optimizer/schedule, precision, and LoRA configuration (outlined in Table 8); only the underlying architecture varies unless otherwise noted.
The implementational details of Section 4.5 is provided in Appendix M.
@@ G.2 Implementation Details and Shared Hyperparameters
G.3 Model Size and Computational Budget
We evaluate a primary backbone model and study transfer across model sizes using a shared training/adaptation recipe (Appendix G.2). Concretely, our main experiments use Llama-2-7B ($\approx$7B parameters). For transferability, we apply coresets selected with a Llama-2-7B scorer to Llama-2-13B ($\approx$13B parameters) and Mistral-7B (v0.3) ($\approx$7B parameters). For completeness, Appendix N additionally reports earlier experiments using Llama-3.2-1B ($\approx$1B parameters) and Llama-3.1-8B ($\approx$8B parameters).
All runs were executed on a single NVIDIA H200 GPU on an internal cluster (CUDA 12.1, PyTorch 2.1, HuggingFace Transformers 4.46). We report wall-clock time or GPU-hours in Section 5.
Appendix H Hyperparameter Ablations on MMLU
We ablate key design choices in TRIM on MMLU while holding the rest of the selection and fine-tuning pipeline fixed to the main setup. Unless stated otherwise, each configuration selects a 5% coreset from the same candidate pool and fine-tunes the same base model as in the corresponding MMLU experiment. We report MMLU test accuracy (%; mean $\pm$ std. over 3 seeds).
Default configuration.
Unless otherwise noted, we use the following TRIM configuration for MMLU: saliency aggregated over the last L=6L{=}6 layers; robust pooling with $0.5\cdot\textsc{Mean}+0.5\cdot\textsc{Max}+0.05\cdot\textsc{Coverage}$ (Appendix C); no IDF reweighting; an NFF penalty of $\lambda{=}0.9$; and scoring using only the last checkpoint. Default values are highlighted in each table.
H.1 Number of layers used for saliency aggregation
In Table 9, we vary the number of final transformer layers $L$ whose attention-derived saliency signals are aggregated when building token fingerprints.
Table 9: MMLU ablation: number of last layers $L$ used for saliency aggregation. Default value highlighted.
# Layers ($L$) MMLU Accuracy (%)
Last 33 48.78 (0.3)
@@ H.1 Number of layers used for saliency aggregation
Last 1212 49.43 (0.2)
H.2 Pooling function for candidate scoring
In Table 10, we compare pooling choices used to aggregate token-level matches into a sample-level score, including mean pooling, max pooling, and coverage-based pooling (Appendix C).
Table 10: MMLU ablation: pooling function used for candidate scoring. Default value highlighted.
Pooling MMLU Accuracy (%)
Mean only 48.26 (0.1)
Max only 49.14 (0.3)
Mean + $0.05\cdotCoverage$ 48.75 (0.3)
$\cdotMean + 0.5+\,0.5\cdotMax$ 48.75 (0.1)
Robust: $\cdotMean + 0.5+\,0.5\cdotMax$ + $0.10\cdotCoverage$ 48.77 (0.2)
@@ H.2 Pooling function for candidate scoring
H.3 IDF reweighting
In Table 11, we evaluate whether inverse document frequency (IDF) reweighting improves robustness by down-weighting frequent tokens and emphasizing rarer, potentially more informative tokens.
Table 11: MMLU ablation: effect of IDF reweighting. Default value highlighted.
Setting MMLU Accuracy (%)
@@ H.3 IDF reweighting
IDF (on) 49.31 (0.1)
H.4 Penalizing non-fingerprinted similarities
In Table 12, we ablate the fallback penalty $\lambda$ in Equation 12, which controls how strongly we penalize matches involving token classes that are not observed in the target validation set and must be mapped to a nearest fingerprinted class.
Table 12: MMLU ablation: effect of the NFF penalty $\lambda$. Default value highlighted.
$\lambda$ MMLU Accuracy (%)
No NFF 44.79 (0.45)
0.5 46.04 (0.8)
[…]
1.00 48.58 (0.5)
H.5 Checkpoints used for scoring
In Table 13, we ablate the number of checkpoints used during scoring. When using multiple checkpoints, we aggregate per-checkpoint scores using learning-rate scaling, following the convention in (Xia et al., 2024).
Table 13: MMLU ablation: checkpoints used in scoring. Default value highlighted.
Checkpoint setting MMLU Accuracy (%)
Pretrained model 44.26 (0.1)
Earliest warmup checkpoint 46.37 (0.2)
[…]
All warmup checkpoints 49.43 (0.5)
Appendix I Bias and Domain-Shift Robustness of Attention-Derived Saliency
To probe whether attention-derived saliency overemphasizes stylistic or high-frequency tokens in a way that harms generalization, we evaluate models fine-tuned on 5% TRIM-selected coresets (targeted to GSM8K) on additional benchmarks beyond GSM8K. Table 14 reports results on MMLU and TruthfulQA (MC2).
Table 14: Out-of-task evaluation on MMLU and TruthfulQA (MC2) for Llama-2-7B. Coresets are 5% selected from the Dolly/Flan_v2/CoT/Oasst1 pool using GSM8K targets (10 examples). Mean accuracy (%) with std. over 3 runs.
Fine-tuning setting MMLU TruthfulQA (MC2)
[…]
LESS 5% coreset 46.30 (0.4) 39.91 (0.2)
[…]
Importantly, for MMLU and TruthfulQA all methods still fine-tune on 5% drawn from the same underlying instruction pool; GSM8K supervision only determines which 5% is selected, not the candidate distribution itself. Accordingly, a GSM8K-targeted 5% coreset need not be optimal for MMLU/TruthfulQA, but it should not systematically degrade performance if the selection procedure introduces no harmful biases. Consistent with this expectation, TRIM-selected coresets do not reduce performance relative to the pretrained model and remain comparable to LESS on both benchmarks, while modestly improving over the pretrained baseline.
Appendix J Robustness to Target Validation Sets
Setup.
All experiments in this section use the same selection setup: we select 5% coresets from the Dolly/Flan_v2/CoT/Oasst1 candidate pool using GSM8K target validation examples, fine-tune Llama-2-7B using LoRA, and evaluate on GSM8K under a fixed 8-shot protocol. Results are mean accuracy (%) with std. over 3 seeds.
J.1 Robustness to validation set size
To quantify TRIM’s sensitivity to the size of the target validation set, we vary the number of target examples used to construct token fingerprints. Table 15 reports GSM8K accuracy as the target set size increases.
Table 15: Effect of target validation set size on GSM8K (8-shot). 5% coreset selected from Dolly/Flan_v2/CoT/Oasst1. Mean accuracy (%) with std. over 3 runs. Default value highlighted.
# Target examples GSM8K accuracy (%)
5 27.37 (0.1)
8 28.43 (0.2)
[…]
20 30.25 (0.2)
50 29.84 (0.1)
100 30.18 (0.4)
Performance improves as the validation set grows: increasing the number of target examples exposes TRIM to a broader set of token contexts, yielding more informative fingerprints. Beyond roughly 20–50 examples, gains saturate and fluctuations remain within the reported standard deviations, indicating that TRIM does not require large target sets to be effective. Even 100 target examples remain negligible relative to the candidate pool, so the added compute and memory overhead is minor in practice.
J.2 Robustness to noisy targets
We further evaluate robustness by injecting noise into the target examples. Specifically, for a randomly chosen fraction of target examples, we swap their answer prompts with those of other validation examples. Table 16 reports GSM8K performance as the noise ratio increases.
Table 16: Robustness to noisy targets on GSM8K (8-shot). Targets: 10 examples. We vary the fraction of noisy target samples used to construct fingerprints. Mean accuracy (%) with std. over 3 runs. Default value highlighted.
Noisy percentage (%) GSM8K accuracy (%)
[…]
10 29.65 (0.1)
25 29.45 (0.4)
50 28.51 (0.1)
75 27.80 (0.1)
100 27.29 (0.1)
Accuracy declines gradually with increasing noise, but the degradation is not catastrophic: even at 100% noise, performance remains comparable to the small-target regime in Table 15. Intuitively, noisy targets still induce attention over similar token support (tokens remain drawn from the same underlying pool), so aggregate token statistics are only partially perturbed. The primary effect of noise is to distort the relative weighting between the query- and key-based saliency components (our $Q/K$ views): mismatched question–answer pairs lead to less task-aligned attention patterns, and the combined $w_{Q}$$w_{K}$ aggregation becomes less informative. Nevertheless, because TRIM aggregates saliency across many tokens and layers, the resulting fingerprints remain reasonably stable, leading to only moderate performance loss even under substantial noise.
Appendix K Source-Subset Composition of Selected Coresets
Figure 4 summarizes how different selection methods draw from the mixed instruction pool (CoT, Flan_v2, Oasst1, Dolly) when constructing 5% coresets for MMLU, BBH, TyDiQA, and GSM8K. The methods exhibit markedly different preferences over sources, but these preferences do not translate into a simple “more from subset $\Rightarrow$ better performance” relationship. This indicates that all sources contain valuable examples and that the main challenge is identifying the most task-relevant instances.
A clear distinction emerges between task-agnostic and task-adaptive selection. TAGCOS produces an almost unchanged mixture across targets, consistent with its task-agnostic design. In contrast, TRIM and LESS adapt their selection distributions to the target task, but they do so using different signals: TRIM relies on token-level representational matching, whereas LESS operates at the sequence level using per-example training dynamics.
Qualitatively, TRIM tends to emphasize reasoning-centric sources when the target requires explicit multi-step inference (e.g., BBH and GSM8K), while leaning more on broad instruction-following/knowledge-heavy sources for MMLU. For TyDiQA, TRIM adopts a more mixed allocation, consistent with multilingual QA benefiting from diversity in instruction style and linguistic coverage rather than a single dominant format. LESS also exhibits target-dependent shifts, but as a sequence-level method it is sensitive to whole-example training signals (rather than localized token overlaps), which can induce different subset preferences even when both methods are task-adaptive.
Appendix L Qualitative Inspection of Token Fingerprints and Nearest-Fingerprinted Fallback (NFF)
What is shown.
We present qualitative examples that make TRIM concrete: (i) representative fingerprinted tokens (token classes that appear in the target validation set and thus receive fingerprints) in Table 17, and (ii) examples of Nearest-Fingerprinted Fallback (NFF), where a token without a fingerprint is mapped to the nearest fingerprinted token in input-embedding space and scored using the mapped fingerprint (with an NFF penalty factor, when enabled) in Table 18.
Table 17: Example fingerprinted tokens (decoded strings) for each target task.
Target task Fingerprinted token (decoded)
MMLU "solution"
"intelligence"
"African"
BBH "Western"
"Furthermore"
"highest"
TyDiQA "ju"
"‘ds‘"
"‘ya‘"
GSM8K "+"
"half"
"Since"
Table 18: Nearest-Fingerprinted Fallback (NFF) examples. For a token without a fingerprint, NFF maps it to the nearest fingerprinted token in input-embedding space and reuses the mapped fingerprint for scoring (optionally down-weighted by the NFF penalty).
Token w/o fingerprint NFF mapped fingerprint token
"solutions" "solution"
"location" "sites"
"players" "games"
"may" "might"
"costs" "bill"
"paid" "bought"
Takeaway.
Across target tasks, the fingerprinted token set contains task-relevant lexical and formatting patterns under the chosen token scope. When a token does not receive a fingerprint, Nearest-Fingerprinted Fallback (NFF) provides a simple backoff by mapping to a nearby fingerprinted token in embedding space.
Appendix M In-Domain Selection with No Target Samples
Setting overview.
We study in-domain selection when there is no external target dataset: the candidate pool itself defines the domain, and the goal is to select a coreset that preserves the pool distribution. We follow the specialized math selection setting of S2L (Yang et al., 2024) using MathInstruct (Yue et al., 2024), a 262K-example instruction-tuning corpus aggregated from 13 math sources. To enable validation-driven selectors without introducing a separate target dataset, we construct a small in-domain target pool $\mathcal{T}_{\text{val}}$ via stratified sampling across sources and use it consistently for all methods that require target/validation examples (e.g., TRIM, LESS). We evaluate this setup with the MAmmoTH/MathInstruct evaluation pipeline and report exact-match accuracy, consistent with S2L.
Data format and training objective.
MathInstruct unifies heterogeneous math datasets into an follow an Alpaca-style instruction-following format (Chen et al., 2024b), where each example is an instruction (with optional input) paired with a response. For supervised fine-tuning, we use standard next-token prediction with the loss masked on the instruction portion, training the model to predict only the response tokens.
Reference target pool for fingerprinting ($\mathcal{T}_{\text{val}}$).
Let the 13 MathInstruct sources be $\{s_{k}\}_{k=1}^{13}$. We build $\mathcal{T}_{\text{val}}$ by sampling $m$ examples uniformly at random from each source, yielding $|\mathcal{T}_{\text{val}}|=13m$. Unless otherwise stated, we use m=20m{=}20 (i.e., $|\mathcal{T}_{\text{val}}|{=}260$). As stated before, $\mathcal{T}_{\text{val}}\subset\mathcal{S}$ by construction. Results are averaged over multiple runs, indicating that a randomly selected $\mathcal{T}_{\text{val}}$ is sufficient for identifying high impact data samples.
Evaluation protocol and benchmarks.
We evaluate on the same six open-form math benchmarks used by S2L under the 0-shot MAmmoTH evaluation pipeline: GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), NumGLUE (Mishra et al., 2022), SVAMP (Patel et al., 2021), Mathematics (Davies et al., 2021), and SimulEq (Koncel-Kedziorski et al., 2016).
Evaluation uses a PoT $\rightarrow$ CoT fallback prompting scheme (attempt program generation/execution first; otherwise fall back to “think step-by-step”) and reports exact match accuracy. Some benchmarks have source overlap with MathInstruct (e.g., GSM8K/MATH/NumGLUE), while others have no source overlap (e.g., SVAMP/Mathematics/SimulEq).
Scoring and training pipeline.
Across all methods, the evaluation model is a pretrained Llama-2-7B that is fine-tuned on the selected subset using the same fixed-epochs protocol. Selection scores are computed using one of two scorer recipes: (i) a warm-started Llama-2-7B scorer (our main-paper default), or (ii) an S2L-style fully-trained proxy scorer (Pythia-70M). We summarize the S2L-style proxy training hyperparameters in Table 19.
Table 19: S2L-style proxy training hyperparameters on MathInstruct (used when training the Pythia proxy to collect loss trajectories).
Proxy model Pythia-70M
Learning rate $2\times 10^{-5}$
Batch size (global) 128
Max sequence length 512
LR schedule / warmup cosine / 3%
Performance comparison.
Table 20 compares the performance of a Llama-2-7B model fine-tuned on $50\%$ coresets selected by various methods using the warm-started Llama-2-7B scorer.
Table 20: Extended evaluation on six open-form math benchmarks for Llama-2-7B fine-tuned on $50\%$ coresets (mean over 3 seeds; std. in parentheses).
Coreset Method Benchmarks with Source Overlap Benchmarks with No Source Overlap
GSM8K MATH NumGLUE Avg SVAMP Mathematics SimulEq Avg
@@ Appendix I Bias and Domain-Shift Robustness of Attention-Derived Saliency
Random 48.16 (0.5) 18.22 (0.3) 55.12 (0.4) 40.50 63.30 (1.2) 38.75 (0.2) 45.16 (0.3) 49.07
LESS 52.10 (0.1) 32.20 (0.1) 60.30 (0.2) 48.20 65.45 (0.2) 41.80 (0.1) 48.75 (0.1) 52.00
S2L 52.45 (0.1) 32.10 (0.2) 60.10 (0.1) 48.22 65.30 (0.2) 41.92 (0.3) 49.26 (0.1) 52.16
TAGCOS 50.20 (0.2) 28.28 (0.1) 58.31 (0.6) 45.60 64.15 (0.4) 40.84 (0.1) 48.85 (0.2) 51.28
[…]
Ablation on target-pool size.
Table 21 varies the number of target samples $m$ per source used to construct $\mathcal{T}_{\text{val}}$, holding all other components fixed.
Table 21: Ablation on the size of the target pool $\mathcal{T}_{\text{val}}$ used for fingerprinting/validation. We sample $m\in\{1,4,10,20\}$ examples per source (13 sources), so $|\mathcal{T}_{\text{val}}|=13m$. We report exact-match accuracy (%) on six open-form math benchmarks (mean over 3 seeds; std. in parentheses).
$m$ per source $|\mathcal{T}_{\text{val}}|$ Benchmarks with Source Overlap Benchmarks with No Source Overlap
GSM8K MATH NumGLUE Avg SVAMP Mathematics SimulEq Avg
1 13 46.20 (0.3) 17.40 (0.1) 52.12 (0.3) 38.57 62.14 (0.4) 39.23 (0.1) 44.00 (0.4) 48.46
4 52 49.51 (0.1) 24.20 (0.3) 55.47 (0.4) 43.06 62.43 (0.2) 40.41 (0.2) 45.16 (0.1) 49.33
10 130 51.89 (0.4) 30.10 (0.6) 59.18 (0.3) 47.06 65.42 (0.2) 40.57 (0.1) 47.25 (0.1) 51.08
[…]
Ablation on scorer recipe.
Finally, Table 22 compares selection scores computed using (i) the warm-started Llama-2-7B scorer and (ii) the fully-trained Pythia-70M proxy, while fine-tuning the same pretrained Llama-2-7B evaluation model under the fixed-epochs protocol.
Table 22: Ablation on the scorer/reference-model recipe used to compute selection scores. Both settings use the same $\mathcal{T}_{\text{val}}$ construction (default m=20m{=}20 per source) and fine-tune the same pretrained Llama-2-7B evaluation model under the fixed-epochs protocol (mean over 3 seeds; std. in parentheses).
Selector recipe for scoring Benchmarks with Source Overlap Benchmarks with No Source Overlap
GSM8K MATH NumGLUE Avg SVAMP Mathematics SimulEq Avg
Warm-start Llama-2-7B 52.23 (0.2) 31.40 (0.3) 60.52 (0.1) 48.05 65.12 (0.4) 42.13 (0.4) 48.75 (0.2) 52.00
Fully-trained Pythia-70M 53.40 (0.1) 30.89 (0.1) 60.46 (0.3) 48.25 64.89 (0.2) 41.78 (0.2) 47.50 (0.2) 51.39
Appendix N Supplementary Experiments on Llama-3.2-1B
Scope.
In addition to the main-paper experiments (which use Llama-2-7B as the primary backbone and study transfer to Llama-2-13B and Mistral-7B (v0.3)), we report here supplementary results from an earlier evaluation suite using Llama-3.2-1B. These results follow the same warmup $\rightarrow$ score $\rightarrow$ select $\rightarrow$ finetune pipeline and are included to demonstrate that the trends observed in the main paper are consistent across model families and benchmark choices.
N.1 Setup: Commonsense Multiple-Choice Benchmarks
We evaluate on three standard commonsense multiple-choice benchmarks: CommonsenseQA (Talmor et al., 2019), SocialIQA (Sap et al., 2019), and HellaSwag (Zellers et al., 2019). We use the same OpenAI-style instruction-tuning evaluation protocol (Wang et al., 2023) and evaluate using the EleutherAI lm-evaluation-harness (Sutawika et al., 2023). We report accuracy on all tasks. Selection uses a small target validation set $T_{\text{val}}$, and evaluation uses fixed in-context exemplars with shot counts consistent with Table 7; evaluation exemplars are drawn from development pools disjoint from $T_{\text{val}}$. All prompts and decoding settings are held fixed across methods.
N.2 Budgeted Accuracy on Llama-3.2-1B
Baselines.
We compare against the same families of coreset methods as in the main paper (Appendix D), and all methods use the standardized warmup-and-selection pipeline (Appendix G).
Coreset Method CommonsenseQA SocialIQA HellaSwag Average Mean
@@
Random 34.05 (1.2) 43.84 (0.1) 48.21 (0.1) 42.03
DSIR (Xie et al., 2023) 37.16 (0.7) 44.38 (0.5) 48.75 (0.2) 43.43
S2L (Yang et al., 2024) 34.10 (0.1) 43.32 (0.2) 48.57 (0.2) 41.99
RDS (Hanawa et al., 2021) 36.16 (1.1) 43.77 (0.1) 48.43 (0.3) 42.79
CLD (Nagaraj et al., 2025) 33.10 (0.4) 43.25 (0.5) 48.35 (0.1) 41.56
LESS (Xia et al., 2024) 39.10 (0.7) 44.52 (0.3) 49.01 (0.1) 44.21
TAGCOS (Zhang et al., 2025b) 34.72 (0.8) 43.70 (0.2) 48.76 (0.1) 42.39
@@
Figure 5: TRIM vs. top baselines on Llama-3.2-1B across coreset budgets.
Results and Takeaways.
As summarized in Table 23, TRIM achieves the strongest overall performance under a fixed 5% budget, matching the main-paper trend that token-centric selection yields more accurate coresets than sample-level baselines. Moreover, Figure 5 shows that TRIM’s advantage is stable across coreset sizes, indicating robustness to budget choice and supporting its use as a reliable selector under varying compute constraints.
N.3 Transferability from a Small Scorer
We assess whether a small scorer can curate data that transfers to larger models. Using Llama-3.2-1B as the scorer, we select a single 5% coreset and reuse it to fine-tune two larger target models: Llama-3.1-8B and Mistral-7B (v0.3). We compare (i) TRIM-Transfer, the coreset chosen by the 1B scorer, and (ii) TRIM-Oracle, a coreset selected using the target model itself as the scorer.
Llama-3.1-8B Mistral-7B (v0.3)
Coreset Method CSQA SIQA HS Avg. CSQA SIQA HS Avg.
[…]
Results and Takeaways.
Table 24 shows that coresets selected by the Llama-3.2-1B scorer transfer with high fidelity to both a larger in-family model and an out-of-family target. Consistent with the main-paper findings, TRIM-Transfer closely tracks (and in some cases matches) the TRIM-Oracle reference, suggesting that TRIM’s token-level fingerprints capture signals that generalizes across model scale and architecture.
N.4 Subset Distribution of Selected Examples
Figure 6: Distribution of selected examples across training subsets for different methods and target tasks in the Llama-3 supplementary setup. Each bar shows the percentage breakdown of selected data from CoT, Dolly, Flan_v2, and Oasst1.
We present the distribution of selected examples across different data selection methods under the Llama-3 evaluation setup in Figure 6.
TAGCOS, being a task-agnostic method, maintains nearly identical selection distributions across all four tasks in this setting (approximately 38.5% CoT, 35.0% Flan_v2, 20.7% Oasst1, and 5.8% Dolly). In contrast, both TRIM and LESS demonstrate task-adaptive behavior, substantially varying their selection patterns based on the specific characteristics of each target task. In our observations for the Llama-3 setup, we find that TRIM predominantly selects data from the Flan_v2 dataset for both SocialIQA (81.3%) and CommonsenseQA (71.8%). This preference is intuitive, as both tasks involve question answering with multiple-choice or short-answer formats, which are well represented in Flan_v2’s diverse instruction-following data. For HellaSwag, TRIM adopts a more balanced approach, selecting comparable amounts from Flan_v2 (56.0%) and CoT (42.7%), consistent with HellaSwag’s commonsense inference format, which benefits from both broad instruction-following and explicit reasoning demonstrations.